OpenStudy (donnie1999):

can someone graph 5x-3y=15 please will medal

OpenStudy (fanduekisses):

Ok so first you need it in slope form: y= mx + b

OpenStudy (fanduekisses):

m is the slope and b is the y-intercept, that is why that form is better for graphing. You plot the y-intercept then use slope to go either up or down and then to the left.

OpenStudy (fanduekisses):

But get it in the y= mx + b form first, do this by getting y by itself so subtract 5x to both sides first.

OpenStudy (fanduekisses):

you'll get -3y= -5x+ 15 You still don't have y by itself, there is a -3 in front. Let's get rid of it by doing the opposite, divide -3 to both sides...

OpenStudy (donnie1999):

the slope is 5/3

OpenStudy (fanduekisses):

We get y=5/3y-5

OpenStudy (fanduekisses):

Now you can graph, start at y= -5, (0,-5)

OpenStudy (fanduekisses):

Use slope, which is 5/3, from -5 go up 5 and 3 to the right.

OpenStudy (donnie1999):

yah I did that so will the next point be 0,3 ? and do I have to shade anything?

OpenStudy (fanduekisses):

Yes you got it but you don't need to shade anything, unless it were an inequality.

OpenStudy (donnie1999):

oh ok thanks :D

OpenStudy (fanduekisses):

You should then end up with a line, just connect (0,-5) with (0,3)

OpenStudy (fanduekisses):

You welcome! ^_^